Wine and Roses, Inc. offers a 7% coupon bond with semiannual payments and a yield to maturity of 7.73%. The bonds mature in 10 years. What is the market price of a $1,000 face value bond?
Answer #1
Current price of bonds = =-pv(rate,nper,pmt,fv)
= $ 949.80
Where,
rate 7.73%*6/12 = 0.03865
nper 10*12/6 = 20
pmt 1000*7%*6/12 = $       35.00
fv = $ 1,000.00